OpenStudy (anonymous):

Determine the volume of the solid of revolution formed by the area enclosed by the curve

OpenStudy (anonymous):

\[\left(\begin{matrix}y ^{2} \\ 16\end{matrix}\right)=1-\left(\begin{matrix}x ^{2} \\ 4\end{matrix}\right)\] (x axis) x=-2 x=2 ordinates